OpenStudy (mantar0078):

PLEASE HELP Identify the asymptotes, domain, and range of the function. Check all that apply. f(x)=1/(x+4)+3 asymptotes: x = -4, y = 3 R:{y|y≠3} asymptotes: x = 3, y = -4 D:{x|x≠−4} R:{y|y≠−4} D:{x|x≠3}

OpenStudy (solomonzelman):

\(\displaystyle f(x)=\frac{1}{x+11}-8\) (1) vertical shifts \(8\) units down, (2) vertical asymptote at \(x=-11\), and (3) horizontal asymptote at \(x=-8\).

OpenStudy (solomonzelman):

Why? I assume you know (1)! Reason (2) \(x+11\ne0\), because otherwise you are dividing by zero. So, \(\displaystyle \frac{1}{x+11}\ne 0\). Note then, that if \(\displaystyle g(x)=\frac{1}{x+11}\), then you therefore have a horizontal asymptote at \(x=0\) (because \(g(x)\ne0\), however the closer you tend to zero, the bigger value you get, going into \(\pm\) infinities on both sides of \(y=0\). Then, \(\displaystyle f(x)=\frac{1}{x+11}-8\) is same, but all of this reasoning is shifted to \(y=-8\).

OpenStudy (solomonzelman):

and vertical asymptote is relevantly simple. if not the concept, for the very least you know that \(x=-11\) gives you a vertical asymptote, because if \(x=-11\) then \(x+11=0\) which is not allowed.

OpenStudy (solomonzelman):

(not allowed due to the fact that you are dividing by 0)

OpenStudy (mantar0078):

How am I supposed to find what y=?

OpenStudy (solomonzelman):

how are you supposed to find the horizontal asymptote?

OpenStudy (mantar0078):

Yes?

OpenStudy (mantar0078):

I've never heard that term before

OpenStudy (solomonzelman):

OK, fine .... :)

OpenStudy (solomonzelman):

how about "vertical asymptote" ?

OpenStudy (mantar0078):

I've heard that one

OpenStudy (mantar0078):

The vertical asymptote would be -4?

OpenStudy (solomonzelman):

We will plot a function \(\displaystyle f(x)=\frac{1}{x-2}+5\) https://www.desmos.com/calculator/ve1u34ikjs

OpenStudy (solomonzelman):

Let's make a flashback.

OpenStudy (solomonzelman):

What is smaller 1/100 or 1/9999 ? (And why?)

OpenStudy (mantar0078):

1/9999 because it's a smaller odd

OpenStudy (solomonzelman):

WRONG !

OpenStudy (solomonzelman):

1/9999 is smaller because you are dividing by a larger number.

OpenStudy (solomonzelman):

that is the way to think about it.

OpenStudy (solomonzelman):

Right? The larger number you divide by, the smaller the result. and the smaller number you divide by, the bigger the result. AGREE ?

OpenStudy (mantar0078):

Yes, I agree. 1/100 would be a smaller decimal than 1/9999

OpenStudy (mantar0078):

Sorry, larger

OpenStudy (solomonzelman):

I was almost despaired before you fixed that :)

OpenStudy (mantar0078):

Haha

OpenStudy (solomonzelman):

OK, so here we come to the reason why division by 0 is invalid ...

OpenStudy (solomonzelman):

1/0.1=10 1/0.05=20 1/0.009=1250 ....

OpenStudy (solomonzelman):

And this is making sense, right (?) Because the smaller the number we divide by, the larger the result, right?

OpenStudy (solomonzelman):

Note as well, that you can choose a number that is very very close to zero, and there are infinitely many (real) numbers that are going closer and closer to zero, right?

OpenStudy (mantar0078):

Yes

OpenStudy (mantar0078):

It's making sense

OpenStudy (solomonzelman):

So, the closer x is to 0, the bigger the quotient 1/x is.

OpenStudy (solomonzelman):

Well, for any number that is very close to zero you may think, I can come up with the number that is closer to zero. (For the very least part, this is true, because if your number is \(x\), then \(y=x/2\) is closer to zero.)

OpenStudy (solomonzelman):

So, this x that approaches 0, will lower than anything, and thus (intuitively) the quotient 1/x is going to be bigger than anything (i.e. infinity).

OpenStudy (solomonzelman):

and the logic is the same, for negatives. In other words, if you approach 0 from the negative side (e.g. -0.0004 and smaller values on that are tending to 0)

OpenStudy (solomonzelman):

with an exception that, a very "small" negative decimal (that is smaller than anything) (such as -0.0000000400123) (will call it x) when 1/x, you get negative infinity.

OpenStudy (solomonzelman):